Which expression below produces the nth term in the sequence?
24 , 30 , 36 , 42 , 48 , ...

Answer by palanisamy(496)

The given sequence is
24 , 30 , 36 , 42 , 48 , ...
This an arithmetic sequence.
First term a = 24
common difference d = 6
nth term = a+(n-1)d
= 24+(n-1)6
= 24+6n-6
= 6n+18